OpenStudy (anonymous):

What are imaginary numbers, and how do I do problems with them? Please help! I can't find a simple explanation anywhere!

OpenStudy (anonymous):

The further I move in school, the more I am convinced that I am a complete idiot and question how I even made it this far in the education system. I am HORRIBLE at math. Anyway, what are imaginary numbers? How do I use them? What the hell do I do with them?! I'm simplifying and solving equations at the moment. Here are a few examples on what I'm doing. (-3i)(4i)(-5i) (7-6i)+(9-11i) 5n^2+35=0 15-28i=3l+4(m)i Also... When in the world am I going to use this in real life?!

OpenStudy (anonymous):

Electrical current can have imaginary numbers in the equations sometimes. And differential equations might get them too.

OpenStudy (anonymous):

\[i=\sqrt{-1}\] which is normally impossible to do because \[i^2=-1\] can't be solved.

OpenStudy (anonymous):

that's why it's not a 'real' number, but a complex one.

OpenStudy (anonymous):

you can draw complex number much like a normal x-y graph. Where the y-axis is your imaginary axis and the x-axis is the 'real number' axis.

OpenStudy (anonymous):



OpenStudy (anonymous):

a bit like that, i drawn 5+3i there.